# Gravitational $D$-Form Factor: The $\sigma$-Meson as a Dilaton confronted with Lattice Data

### v1

@@ 4.2 The pion gravitational form factor
$D^{\pi}(q^{2})=\frac{q^{2}\,r_{\mathrm{E},\sigma}^{\pi}}{q^{2}-m_{\mathrm{E},\sigma}^{2}}-1+\hat{b}+b^{{}^{\prime}}q^{2}+\frac{r_{\mathrm{eff}}}{q^{2}-m_{\mathrm{eff}}^{2}}\;,$ (4.6)
@@
which the fits satisfy. An alternative would be to impose the constraint along with next-to-leading order (NLO) computation in chiral perturbation theory [75].
As in the nucleon case, we set $r_{\mathrm{eff}}=0$ in the main fit. The theoretical prediction for the residue is $r_{\sigma}^{\pi}=2/3$ (2.8), with uncertainties arising from NLO chiral corrections of order ${\cal O}([m_{\pi}^{2}\ln m_{\pi}]/(4\pi F_{\pi})^{2})$. In the chiral limit $m_{u,d,s}\to 0$, the value of $F_{\pi}$ decreases by about $10\%$ [91]. Adopting this as a guideline for estimating the residue in the chiral limit, and rounding the uncertainty generously, we arrive at the prediction
@@
Table 2: Same as table 1, but now for the pion data [17] fitted using the parametrisation (4.6). LO dilaton effective theory predicts $r_{\sigma}^{\pi}=2/3$. The correlations between parameters for $m_{\mathrm{E},\sigma}=550\,\mbox{MeV}$ are $(\rho_{r\hat{b}},\,\rho_{rb^{\prime}},\,\rho_{\hat{b}b^{\prime}})=(-0.94,\,0.86,\,-0.65)$.
We apply the same fitting procedure as for the nucleon using all 24 data points of ref. [17], with $N_{\mathrm{dof}}=21$ since there are three fit parameters. The main fits are shown in Fig. 1 and reported in table 2 for a range of effective $\sigma$-masses, with uncertainties represented as in table 1. The dependence on the effective mass is much smaller in this case, and our main fit result is
@@
    $(r_{\mathrm{E},\sigma},\hat{b})^{\pi}$: with $r_{\mathrm{E},\sigma}^{\pi}=0.80(8)$, $\hat{b}=-0.006(62)$ and $\hat{\chi}^{2}=1.30$
@@
    $(r_{\mathrm{E},\sigma},\hat{b},r_{\mathrm{eff}})^{\pi}$: with $r_{\mathrm{E},\sigma}^{\pi}=0.08(33)$, $\hat{b}=0.93(42)$, $r_{\mathrm{eff}}=0.77(34)$ and $\hat{\chi}^{2}=1.12$
@@
    $(r_{\mathrm{E},\sigma},r_{\mathrm{eff}})^{\pi}$: with $r_{\mathrm{E},\sigma}^{\pi}=0.81(4)$, $r_{\mathrm{eff}}=0.02(5)$ and $\hat{\chi}^{2}=1.29$
@@
    $(r_{\mathrm{E},\sigma},b^{\prime},r_{\mathrm{eff}})^{\pi}$: with $r_{\mathrm{E},\sigma}^{\pi}=0.58(13)$, $b^{\prime}=-0.11(6)$, $r_{\mathrm{eff}}=-0.09(8)$ with $\hat{\chi}^{2}=1.19$
@@
The situation is different for case b). Unlike for the nucleon, the extracted residue is not compatible with the dilaton prediction. This indicates that the converse approach, establishing $\sigma$-dominance directly, is not feasible for the pion. Rewriting the result as per footnote 4, one finds a residue of $m_{\sigma}^{2}\approx 0.3\,\mbox{GeV}^{2}$, far smaller than both the nucleon residue (4.2) and the potential $f_{0}(980)$ contribution in the pion case. This outcome is therefore not surprising and also explains why the fits are less sensitive to the effective mass $m_{\mathrm{E},\sigma}$ than in the nucleon case.
@@
4.3 A quick note on the $D$-form factor in the infrared - the $D$-term
As previously discussed, conserved-current form factors are often connected with simple physical interpretations in the infrared. For instance, the gravitational form factors satisfy $A(0)=1$ and $J(0)=\frac{1}{2}$, reflecting their association with energy and angular momentum. By contrast, an interpretation of the $D$-form factor has long been elusive; for reviews, see [3, 6]. Particular emphasis has been given to $D\equiv D(0)$, commonly referred to as the $D$-term (or Druck-term). Viewing the $\sigma$-meson as a dilaton, combined with phenomenological fits, provides a new perspective on this longstanding problem.
@@ 4.3 A quick note on the $D$-form factor in the infrared - the $D$-term
as emphasised earlier. While this scenario is generally not considered likely, there is also no definitive evidence ruling it out, and recent $N_{f}=4$ quark-mass-degenerate lattice simulations do certainly not exclude a massless $\sigma$ [55]. [^6] See also [94] for functional methods or holographic approaches [95].
In the real world, where the $\sigma$ acquires a nonzero mass, at least due to light quarks, the pole disappears and the $D$-term remains finite. In the Breit frame, the $D$-term can be expressed in terms of the pressure and shear forces [40]. When supplemented with the assumption of mechanical stability [96], this formulation implies $D<0$ (or even D<0.20(2)D<-0.20(2) [97]), a feature observed in many hadronic systems though not, for example, in hydrogen [98, 99, 100]; see [6] for a thorough review. In our approach we obtain (3.8),
@@
Other values obtained from the lattice data include the $z$-expansion fit D=3.35(58)D=-3.35(58) in the original paper [16] and D=3.0(4)D=-3.0(4) from a constrained tripole-fit [29]. Furthermore, a dispersive analysis at physical pions mass was used to obtain a value of $D=-3.38^{+0.34}_{-0.35}$ [28]. The close agreement among these results is reassuring. However, our main point is not the value but that the $\sigma$-contribution is necessarily negative in the dilaton picture, and if dominant over the background, it implies a negative $D$-term. It is in this respect that our work differs from others.
The $D$-term is also sought after since it enter the radii associated with the energy-momentum tensor. The mean square radius in the Breit-frame and the mass radius read [3]
@@
respectively. From (4.11) we infer that the $\sigma$-contribution (being nonchalent, in distinguishing $m_{N}$ and $\bar{m}_{N}$) is inversely proportional to the effective $\sigma$-mass
$\langle r^{2}_{\Theta}\rangle|_{\sigma}=6(A^{\prime}(0)+\frac{1}{m_{\mathrm{E},\sigma}^{2}}-\frac{3}{4}\frac{b(0)}{m_{N}^{2}})\;.$ (4.14)
Thus the mass radii gets larger as the effective $\sigma$-mass decreases. This is intuitive and analogous to the pion charge radius, although there the dependence is only logarithmic on the pion mass [67].
What about the $D$-term for other hadrons? For non-Goldstone states the same pattern applies, see (1.2). In particular, we find similar behaviour for the gluonic gravitational form factors of the $\rho$-meson and the $\Delta$-baryon [101], based on lattice data at $m_{\pi}\approx 450\,\mbox{MeV}$ [15]. Let us turn to the Goldstone case. For the pion, the soft-pion theorem requires $D^{\pi}(0)=-1$ [75]. However, in the presence of a massless $\sigma$, this relation is modified to $D^{\pi}(0)=-\tfrac{1}{3}$ [71], as follows directly from (2.8), since the dilaton pole prevents a naive application of the soft-pion theorem. [^7] For a massive $\sigma$, one gets $D^{\sigma}(0)=-\frac{3+2\Delta_{\cal O}}{3}$ with $\Delta_{\cal O}$ the scaling dimension the mass generating operator $\cal O$}. If ${\cal O}=m_{q}\bar{q}q$ then $\Delta_{\cal O}=2$ and $D^{\sigma}(0)=-\frac{7}{3}$. For a massless $\sigma$, one has $D^{\sigma}(0)=-\frac{1}{3}$ as for the pion [71].
5 Conclusions and Discussion
In this work, we have fitted the pion and the nucleon gravitational form factors, from lattice QCD data [16, 17], to a parametrisation (3.8) where the $\sigma$-meson plays a central role. The main quantity of interest is the residue $r_{\mathrm{E},\sigma}$ of the effective $\sigma$-pole which we compare to the dilaton effective theory predictions (2.8). The fit ansatz was motivated theoretically and tested specifically in the linear $\sigma$-model which served as a toy model, allowing us to understand the difference between the complex-valued and the effective Euclidean residue.
@@ 5 Conclusions and Discussion
the best fit and the dilaton prediction, which are compatible with each other. The first uncertainty arises from the lattice data, and the second estimates the model-uncertainty by varying the effective $\sigma$-mass. Additionally, different background parametrisations confirm the robustness of this result, and although extra terms can partially mimic the role of the $\sigma$-pole, the extracted residue remains stable within uncertainties.
@@
the best fit and the dilaton prediction. The uncertainties are of the same type as for the nucleon. Importantly, while the data are consistent with a dilaton-like residue, the converse, establishing $\sigma$-dominance directly from the lattice results, is not feasible. This might well be due to the data and the somewhat small residue of the pion as a Goldstone boson, and does not undermine the overall consistency of the dilaton interpretation.
Overall, our findings support the idea that QCD is governed by an infrared fixed point with the $\sigma$-meson becoming a light or massless dilaton (in the limit of vanishing light quark masses). Presumably, the $\sigma$-meson plays a key role in understanding the gravitational $D$-form factor. In its purest form, with a massless $\sigma$, the nucleon form factor develops a pole in the infrared $D^{N}(q^{2})=\frac{4}{3}\frac{m_{N}^{2}}{q^{2}}+{\cal O}(1)$. For finite $\sigma$-mass, relevant to the real world with nonzero quark mass, our analysis in section 4.3 implies that $D^{N}(0)<0$ holds, provided that the $\sigma$-term dominates.
As seen from Eqs. (5.1) and (5.2), the dominant uncertainties stem from the data. Further progress can be achieved by improving the precision of the data, extending the kinematic $q^{2}$-range and by going to lower quark masses. Other potential directions, are simulating with degenerate quarks (as in beyond-the-Standard-Model studies) or dispersive approaches [28, 30]. In the longer term we might hope for competitive information from experiment, including the future electron-ion collider [102]. Another avenue is testing other systems, varying the spin, for which $\rho$-meson and the $\Delta$-baryon gluonic gravitational form factors are available at $m_{\pi}\approx 450\,\mbox{MeV}$ [15]. This comes with its very own set of challenges, partly related to the proton mass decomposition [103] and will therefore be discussed in a forthcoming paper [101].
@@ Acknowledgements:
RS and RZ are supported by the STFC via the consolidated grants ST/T000600/1 and ST/X000494/1. We are grateful to Max Hansen, Keh-Fei Liu, Feng-Kun Guo, Mannque Rho, David Schaich and Hanquing Zheng for discussions. A very special thanks go to the MIT-group members Daniel Hackett, Patrick Oare, Phiala Shanahan and especially Dimitra Pefkou for providing us with their form factor data. Analytic computations were performed with FeynCalc [104] and some Passarino-Veltman functions were numerically checked with LoopTools [105] for real kinematics.
Appendix A The Gell-Mann Lévy linear $\sigma$-model – a study case
We use the linear $\sigma$-model [106] as a toy model to illustrate the treatment of particles with a sizeable width, thereby providing a basis for interpreting the fit ansatz of Eq. (3.8). Prior to discussing the underlying physics of the linear $\sigma$-model and the details of the corresponding calculation, we first introduce the relevant nucleon form factor $F(q^{2})$ [^8] We may regard this form factor as a toy model for the improvement term $T_{\mu\nu}\supset\frac{1}{3}(\partial_{\mu}\partial_{\nu}-\eta_{\mu\nu}\partial^{2})\sigma$ (2.4).

### v2

@@ 4.2 The pion gravitational form factor
$D^{\pi}(q^{2})=\frac{q^{2}\,r_{\mathrm{E},\sigma}^{\pi}}{q^{2}-m_{\mathrm{E},\sigma}^{2}}-1+\hat{b}+b^{{}^{\prime}}q^{2}+b^{{}^{\prime\prime}}q^{4}+\frac{r_{\mathrm{eff}}}{q^{2}-m_{\mathrm{eff}}^{2}}\;,$ (4.6)
@@
which the fits satisfy. An alternative would be to impose the constraint along with next-to-leading order (NLO) computation in chiral perturbation theory Donoghue:1991qv .
As in the nucleon case, we set $r_{\mathrm{eff}}=0$ in the main fit. The theoretical prediction for the residue is $r_{\sigma}^{\pi}=2/3$ (2.9) For the uncertainties, we include the $13\%$ $q^{2}$-dependent corrections, as in the nucleon case. Pion mass corrections, ${\cal O}([m_{\pi}^{2}\ln m_{\pi}]/(4\pi F_{\pi})^{2})$, are about four times smaller, so that adding the contributions in quadrature leads to
@@
Table 2: Same as table 1, but now for the pion data Hackett:2023nkr fitted using the parametrisation (4.6). LO dilaton effective theory predicts $r_{\sigma}^{\pi}=2/3$. The correlations between parameters for $m_{\mathrm{E},\sigma}=550\,\mbox{MeV}$ are $(\rho_{r\hat{b}},\,\rho_{rb^{\prime}},\,\rho_{\hat{b}b^{\prime}})=(-0.94,\,0.86,\,-0.65)$.
We apply the same fitting procedure as for the nucleon using all 24 data points of ref. Hackett:2023nkr , with $N_{\mathrm{dof}}=21$ since there are three fit parameters. The main fits are shown in Fig. 1 and reported in table 2 for a range of effective $\sigma$-masses, with uncertainties represented as in table 1. The dependence on the effective mass is much smaller in this case, and our main fit result is
@@
    $\{r_{\mathrm{E},\sigma},\hat{b}\}^{\pi}=\{\,0.80(8)\,,\,-0.006(62)\}$ and $\hat{\chi}^{2}=1.30$
@@
    $\{r_{\mathrm{E},\sigma},\hat{b},r_{\mathrm{eff}}\}^{\pi}=\{\,0.24(27)\,,\,0.81(38)\,,\,0.96(43)\}$ and $\hat{\chi}^{2}=1.13$
@@
    $\{r_{\mathrm{E},\sigma},r_{\mathrm{eff}}\}^{\pi}=\{\,0.81(4)\,,\,0.03(7)\}$ and $\hat{\chi}^{2}=1.29$
@@
    $\{r_{\mathrm{E},\sigma},b^{\prime},r_{\mathrm{eff}}\}^{\pi}=\{\,0.57(14)\,,\,-0.11(6)\,,\,-0.13(11)\}$ and $\hat{\chi}^{2}=1.19$
  • e)
    $\{r_{\mathrm{E},\sigma},\hat{b},b^{\prime},b^{\prime\prime}\}^{\pi}=\{\,0.23(35)\,,\,0.16(11),\,-0.37(29)\,,\;-0.10(11)\}$ and $\hat{\chi}^{2}=1.19$
@@
The situation is different for case b). Unlike for the nucleon, the extracted residue is marginally compatible with the dilaton prediction. This indicates that the converse approach, establishing $\sigma$-dominance directly, is not feasible for the pion. Rewriting the result as per footnote 4, one finds a residue of $r_{\sigma}^{\pi}\to\frac{2}{3}m_{\sigma}^{2}\approx 0.2\,\mbox{GeV}^{2}$ (using $m_{\sigma}\to m_{\mathrm{E},\sigma}$), far smaller than both the nucleon residue (4.2) and the potential $f_{0}(980)$ contribution in the pion case. This outcome is therefore not surprising and also explains why the fits are less sensitive to the effective mass $m_{\mathrm{E},\sigma}$ than in the nucleon case. As for the nucleon, we reduce the fit interval to $0,-1.5]\,\mbox{GeV}^{2}$ and $0,-1]\,\mbox{GeV}^{2}$ for the linear background. For these choices we obtain $\{r_{\mathrm{E},\sigma}^{\pi},\hat{\chi}^{2}\}=\{0.56(19),1.45\}$ and $\{r_{\mathrm{E},\sigma}^{\pi},\hat{\chi}^{2}\}=\{0.32(28),0.46\}$, respectively, which is reasonably stable but at the same time confirms our finding that the pion case is slightly less robust than the nucleon case.
@@
4.3 A quick note on the $D$-form factor in the infrared — the $D$-term
As previously discussed, conserved-current form factors are often connected with simple physical interpretations in the infrared. For instance, the gravitational form factors satisfy $A(0)=1$ and $J(0)=\frac{1}{2}$, reflecting their association with energy and angular momentum. By contrast, an interpretation of the $D$-form factor has long been elusive; for reviews, see Polyakov:2018zvc ; Lorce:2025oot . Particular emphasis has been given to $D\equiv D(0)$, commonly referred to as the $D$-term (or Druck-term). Viewing the $\sigma$-meson as a dilaton, combined with phenomenological fits, provides a new perspective on this longstanding problem.
@@ 4.3 A quick note on the $D$-form factor in the infrared — the $D$-term
as emphasised earlier. While this scenario is generally not considered likely, there is also no definitive evidence ruling it out, and recent $N_{f}=4$ quark-mass-degenerate lattice simulations do certainly not exclude a massless $\sigma$ LatKMI:2025kti . [^7] See also Goertz:2024dnz for functional methods or Elander:2025fpk for holographic approaches.
In the real world, where the $\sigma$ acquires a nonzero mass, at least due to light quarks, the pole disappears and the $D$-term remains finite. In the Breit frame, the $D$-term can be expressed in terms of the pressure and shear forces Polyakov:2002yz . When supplemented with the assumption of mechanical stability Perevalova:2016dln , this formulation implies $D<0$ (or even D<0.20(2)D<-0.20(2) Gegelia:2021wnj ), a feature observed in many hadronic systems though not, for example, in hydrogen Ji:2022exr ; Czarnecki:2023yqd ; Freese:2024rkr ; see Lorce:2025oot for a thorough review. In our approach we obtain (3.8),
@@
Other values obtained from the lattice data include the $z$-expansion fit D=3.35(58)D=-3.35(58) in the original paper Hackett:2023rif and D=3.0(4)D=-3.0(4) from a constrained tripole-fit Broniowski:2025ctl . Furthermore, a dispersive analysis at physical pions mass was used to obtain a value of $D=-3.38^{+0.34}_{-0.35}$ Cao:2024zlf . The close agreement among these results is reassuring. However, our main point is not the value but that the $\sigma$-contribution is necessarily negative in the dilaton picture, and if dominant over the background, it implies a negative $D$-term. It is in this respect that our work differs from others.
The $D$-term is also sought after since it enters the radii associated with the energy-momentum tensor. The mean square radius in the Breit-frame and the mass radius read Polyakov:2018zvc
@@
respectively. From (4.11) we infer that the $\sigma$-contribution is inversely proportional to the effective $\sigma$-mass
$\langle r^{2}_{\Theta}\rangle|_{\sigma}=6(A^{\prime}(0)+\frac{1}{m_{\mathrm{E},\sigma}^{2}}\frac{\bar{m}_{N}^{2}}{m_{N}^{2}}-\frac{3}{4}\frac{b(0)}{m_{N}^{2}})\;.$ (4.14)
Thus the mass radius gets larger as the effective $\sigma$-mass decreases. This is intuitive and analogous to the pion charge radius, although there the dependence is only logarithmic on the pion mass Donoghue:1992dd .
What about the $D$-term for other hadrons? For non-Goldstone states the same pattern applies, see (1.2). In particular, we find similar behaviour for the gluonic gravitational form factors of the $\rho$-meson and the $\Delta$-baryon Stegeman:2025tdl , based on lattice data at $m_{\pi}\approx 450\,\mbox{MeV}$ Pefkou:2021fni . Let us turn to the Goldstone case. For the pion, the soft-pion theorem requires $D^{\pi}(0)=-1$ Donoghue:1991qv . However, in the presence of a massless $\sigma$, this relation is modified to $D^{\pi}(0)=-\tfrac{1}{3}$ Zwicky:2023fay , as follows directly from (2.9), since the dilaton pole prevents a naive application of the soft-pion theorem. [^8] For a massive $\sigma$, one gets $D^{\sigma}(0)=-\frac{3+2\Delta_{\cal O}}{3}$ where $\cal O$} is the operator generating the mass. If ${\cal O}=m_{q}\bar{q}q$ then $\Delta_{\cal O}=2$ and $D^{\sigma}(0)=-\frac{7}{3}$. For a massless $\sigma$, one has $D^{\sigma}(0)=-\frac{1}{3}$ as for the pion Zwicky:2023fay .
5 Conclusion and Discussion
In this work, we have fitted the pion and the nucleon gravitational form factors (from lattice QCD data Hackett:2023rif ; Hackett:2023nkr ) to a parametrisation (3.8) where the $\sigma$-meson plays a central role. The main quantity of interest is the residue $r_{\mathrm{E},\sigma}$ of the effective $\sigma$-pole which we compare to the dilaton effective theory predictions (2.9). The fit ansatz was motivated theoretically and tested specifically in the linear $\sigma$-model which served as a toy model, allowing us to understand the difference between the complex-valued residue and the effective Euclidean residue.
@@ 5 Conclusion and Discussion
the best fit result and the dilaton prediction, which are compatible with each other. The first uncertainty arises from the lattice data, and the second estimates the model-uncertainty by varying the effective $\sigma$-mass. Additionally, different background parametrisations confirm the robustness of this result, and although extra terms can partially mimic the role of the $\sigma$-pole, the extracted residue remains stable within uncertainties.
@@
the best fit result and the dilaton prediction. The uncertainties are of the same type as for the nucleon. Importantly, while the data are consistent with a dilaton-like residue, the converse, establishing $\sigma$-dominance directly from the lattice results, is not feasible. This might well be due to the data and the somewhat small residue of the pion as a Goldstone boson, and does not undermine the overall consistency of the dilaton interpretation.
Overall, our findings support the idea that QCD is governed by an infrared fixed point with the $\sigma$-meson becoming a light or massless dilaton (in the limit of vanishing light quark masses). [^9] If the dilaton Goldberger–Treiman mechanism (2.7) is realised, one can test whether the pion and the sigma decay constants satisfy $F_{\pi}\approx F_{\sigma}$ in the chiral limit. This relation is essential for incorporating a Yukawa-type mechanism into a strongly coupled sector with a dilaton that mimics the behaviour of the Higgs boson Cata:2018wzl ; Zwicky:2023krx . With $\bar{F}_{\pi}=70(3)\,\mbox{MeV}$ in the $SU(3)$ chiral limit Bali:2022qja and assuming $\bar{F}_{\pi}=\bar{F}_{\sigma}$, one finds $g_{\sigma NN}=(\bar{m}_{N}-\Delta m_{N})/\bar{F}_{\sigma}\approx 10$. This falls well within the phenomenological range reported in the literature Wu:2023uva and is therefore deserving of further study, Presumably, the $\sigma$-meson plays a key role in understanding the gravitational $D$-form factor. In its purest form, with a massless $\sigma$, the nucleon form factor develops a pole in the infrared $D^{N}(q^{2})=\frac{4}{3}\frac{m_{N}^{2}}{q^{2}}+{\cal O}(1)$. For finite $\sigma$-mass, relevant to the real world with nonzero quark mass, our analysis in section 4.3 implies that $D^{N}(0)<0$ holds, provided that the $\sigma$-term dominates.
As seen from Eqs. (5.1) and (5.2), the dominant uncertainties stem from the data. Further progress can be achieved by improving the precision of the data, extending the kinematic $q^{2}$-range and by going to lower quark masses. Other potential directions, are simulating with degenerate quarks (as in beyond-the-Standard-Model studies) or dispersive approaches Cao:2024zlf ; Cao:2025dkv . In the longer term we might hope for competitive information from experiment, including the future electron-ion collider Accardi:2012qut . Another avenue is testing other systems, varying the spin, for which $\rho$-meson and the $\Delta$-baryon gluonic gravitational form factors are available at $m_{\pi}\approx 450\,\mbox{MeV}$ Pefkou:2021fni . This comes with its very own set of challenges, partly related to the proton mass decomposition Ji:1994av , the $A$-form factor at zero momentum transfer (extensively studied on the lattice Borsanyi:2020bpd ; Wang:2021vqy ; Liu:2021lke ; Alexandrou:2024zvn ; Alexandrou:2024ozj ), and will therefore be discussed in a forthcoming papers Stegeman:2025tdl ; appear .
@@ Acknowledgements:
[…]
Appendix A The Gell-Mann Lévy linear $\sigma$-model - a Toy Model
We use the linear $\sigma$-model Gell-Mann:1960mvl as a toy model to illustrate the treatment of particles with a sizeable width, thereby providing a basis for interpreting the fit ansatz of Eq. (3.8). Prior to discussing the underlying physics of the linear $\sigma$-model and the details of the corresponding calculation, we first introduce the relevant nucleon form factor $F(q^{2})$ [^10] We may regard this form factor as a toy model for the improvement term $T_{\mu\nu}\supset\frac{1}{3}(\partial_{\mu}\partial_{\nu}-\eta_{\mu\nu}\partial^{2})\sigma$ (2.5).